OpenStudy (anonymous):

The numbers 7, 24, and c are a Pythagorean triple. What is the value of c? A. 31 B. 27 C. 24 D. 25

OpenStudy (calculusxy):

Use the formula a^2+b^2=c^2

OpenStudy (acxbox22):

7^2+24^2=c^2 625=c^2 c=15 D

OpenStudy (calculusxy):

C is the hypotenuse (or the longest side) that you want to solve for by using 7 and 24 as a and b (the legs) .

OpenStudy (anonymous):

25

OpenStudy (anonymous):

is the answer

OpenStudy (calculusxy):

Correct

OpenStudy (anonymous):

I have 3 or 4 more qusetions

OpenStudy (calculusxy):

I'll be happy to help :)